OpenStudy (anonymous):

write 2x+2y=-8 in slope intercept form. My answer y=-x-4??

OpenStudy (anonymous):

U r right.

jhonyy9 (jhonyy9):

2x+2y=-8 divide both sides by 2 x+y=-4 y=-x-4 - yes your answer is right,correct sure